calculate how many grams of sulfur would combine with 26.1 g of potassium to form K2S
1
Expert's answer
2019-03-11T08:18:34-0400

2K+S = K2S

n= m/M

n(K) = 26.1 g/39.1 g/mol = 0.668 mol

As mole ratio of potassium and sulfur is 2:1, then n(S) = n(K)/2 = 0.668 mol/2 = 0.334 mol

m= M*n

m(S)= 32.07 g/mol * 0.334 mol = 10.7 g

Answer: m(S) = 10.7 g